OpenStudy (anonymous):

Find the point-slope form of the equation of the line passing through the points (1,2) and (3,-4). A) y - 2 = -3(x - 1) B) y - 2 = 3(x - 3) C) y + 4 = -3(x - 1) D) y - 2 = 3(x - 1)

OpenStudy (anonymous):

help

OpenStudy (anonymous):

Hey girl

OpenStudy (ankit042):

can you find the slope?

OpenStudy (ankit042):

then use the formula y-y1 = m(x-x1) where m is the slope

OpenStudy (anonymous):

slope= -4-2/3-1 =-6/2=-3...so y-2=-3(x-1)

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

@lilli.tharp

OpenStudy (anonymous):

yes thanks u

OpenStudy (anonymous):

you are welcome:D